OpenStudy (anonymous):

A $4,000.00 principal earns 5% interest, compounded annually. After 4 years, what is the balance in the account? $500,000.00 $4,862.03 $4,600.00 $20,250.00

OpenStudy (anonymous):

A boat costs $92,000.00 and depreciates in value by 15% per year. How much will the boat be worth after 10 years? $18,112.45 $78,200.00 $18,941.98 $69,000.00

OpenStudy (anonymous):

Anyone?

OpenStudy (anonymous):

Free Medal

OpenStudy (anonymous):

@help?!?!?!

OpenStudy (anonymous):

@HailKK

OpenStudy (anonymous):

For the first one I got C and I don't understand the second one

OpenStudy (anonymous):

Give me a few minutes

OpenStudy (anonymous):

92,000 * .15 = x 92,000 - x = value of boat after one year x * .15 = y x - y = value of boat after two years Get it?

OpenStudy (anonymous):

@Anthony2992992 , Do you get it?

OpenStudy (anonymous):

yea but where does y come from?

OpenStudy (anonymous):

y is the same number throughout the whole thing

OpenStudy (anonymous):

y = x * .15

OpenStudy (anonymous):

oh I get it so, y=13800^10x0.15

OpenStudy (anonymous):

wait..... no that's not right

OpenStudy (anonymous):

I don't get it...

OpenStudy (anonymous):

If you meant ^ as a multiplication sign, then you are right

OpenStudy (anonymous):

What grade are you in btw?

OpenStudy (anonymous):

9th

OpenStudy (anonymous):

I was suppose to be in geometry

OpenStudy (anonymous):

You were more likely to get an answer in General Math anyways since there are more people here...

OpenStudy (anonymous):

well is the first one correct?

OpenStudy (anonymous):

Yes...I believe so.

OpenStudy (anonymous):

Are you onlineschooled?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Do you like it?

OpenStudy (anonymous):

Thanks for the medal!

OpenStudy (anonymous):

ya... it not that bad and your welcome